390 - (351 - 536 x (590 - 527)) = ?
1
Expert's answer
2012-07-24T07:22:16-0400
390 - (351 - 536 x (590 - 527)) =390 - (351 - 536 x 63) =390 - 351 + 536 x 63 =39 + 536 x 63 =33807